OpenStudy (anonymous):

Please help....find the equation of a line perpendicular to y=-1/5x+7 and has point (-9,0). I just need a reminder on how to do this.

OpenStudy (compassionate):

Perpendicular lines mean they have opposite slopes. \[y = - \frac{ 1 }{ 5 }x + 7\] \[y = -\frac{ 5 }{ 1 }(-9) + 7 + 0\] I believe that is right? I may be wrong.

OpenStudy (anonymous):

your slope is right, the negative reciprocal of \(-\frac{1}{5}\) is \(5\)

OpenStudy (compassionate):

So it should be, correct? y = 5x + b y = 5(-9) +0 y = -45

OpenStudy (rajee_sam):

Two perpendicular lines with slopes \[m _{1} \space and \space m _{2}\] then \[m _{1} \times m _{2} = -1\]

OpenStudy (compassionate):

@satellite73 I am aware.

OpenStudy (anonymous):

the equation is wrong however you need \(y-y_1=m(x-x_1)\) with \(m=5, x_1=-9,y_1=0\)

OpenStudy (compassionate):

y - y1 = m(x - x1) y - 0 = 5(x - (-9) Right, you're right. It's been awhile since I've done that.

OpenStudy (anonymous):

now you have it \[y=5x+9\]

OpenStudy (rajee_sam):

now the slope of second line will be 5 as you have found out. y= 5x + b now using the point given (-9,0) plug it in the above equation 0 = 5(-9) + b; this will give you b = 45 So the equation of the line is y = 5x + 45

OpenStudy (compassionate):

Last time I did this I was in middle school. Ha, even my memory is blurry.

OpenStudy (rajee_sam):

sorry satellite the answer is \[y = 5x + 45\]

OpenStudy (anonymous):

yes, you are right it is \(y=5x+45\)

OpenStudy (compassionate):

y - y1 = m(x - x1) y - 0 = 5[(x - (-9)] y = 5 (x + 9) y = 5x + 45 Yea, he's right Satelite.

OpenStudy (compassionate):

Don't pick on us Rajee, today isn't our brightest of days when it comes to math. :P

OpenStudy (rajee_sam):

you guys have a 100 and 75 rating I just a mole you guys are mountains

OpenStudy (anonymous):

thank you guys this help me so much i did this just last night and i couldnt remember I just drew a blank. math is not very nice to me :)

OpenStudy (compassionate):

Haha, yea, sometimes even the good mathematicians have brain farts!